Calculate the standard change in Gibbs free energy, Δ?∘rxnΔGrxn∘ , for the given reaction at 25.0 ∘C25.0 ∘C . Consult the table of thermodynamic properties for standard Gibbs free energy of formation values.

KCl(s)↽−−⇀K+(aq)+Cl−(aq)KCl(s)↽−−⇀K+(aq)+Cl−(aq)

Δ?∘rxn=

Determine the concentration of K+(aq)K+(aq) if the change in Gibbs free energy, Δ?rxnΔGrxn , for the reaction is −8.31 kJ/mol−8.31 kJ/mol .

[K+]=
Answer #1

∆G° = ∆G°K+ + ∆G°Cl- - ∆G°KCl = -283.27 -131.23 -(-409.14)

= - 5.36 kJ/mol (Answer)

KCl(s) \rightarrow K+(aq) + Cl-(aq)

[K+] = [Cl-]

∆G = ∆G° + RT ln([K+][Cl-])

-8.31×1000 = -5.36×1000 + 8.314 × 298 × ln[K+]2

ln [K+] = - 0.595

[K+] = e-0.595 = 0.55 M. (Answer)